OpenStudy (anonymous):

write this equation in function form: y - 7 = 1/3(x-9)

OpenStudy (anonymous):

Assuming they want y to be the output of the function, x to be the input, and the function to be called f(x): Add 7 to both sides y = 1/3(x-9) + 7 Replace y with f(x) f(x) = 1/3(x-9) + 7 Any questions?

OpenStudy (anonymous):

sorry. i meant to write in terms of x. i got y = 1/3x +4. is that right?

OpenStudy (anonymous):

Oh, if you simplify the function then yes, it is f(x) = 1/3x - 3 + 7 = 1/3x + 4

OpenStudy (anonymous):

ok. thanks!!!

OpenStudy (anonymous):

Yeah! Would you mind pressing "best response" on one of my responses? Glad I could help!